OpenStudy (anonymous):

sin^4*x+cos^4*x=7:8 x=??

OpenStudy (anonymous):

\[\sin ^{4}x+\cos ^{4}x+2\sin ^{2}x \cos ^{2}x-2\sin ^{2}x \cos ^{2}x=\frac{ 7 }{ 8 }\] \[\left( \sin ^{2}x+\cos ^{2} x\right)^{2}-2\sin ^{2}x \cos ^{2}x=\frac{ 7 }{8}\] \[-2\sin ^{2}x \cos ^{2}x=\frac{ 7 }{ 8 }-1=-\frac{ 1 }{8 }\] \[\left( 2\sin x \cos x \right)^{2}=\frac{ 1 }{ 4 },\sin ^{2}2x=\frac{ 1 }{ 4 },2\sin ^{2}2x=\frac{ 1 }{ 2 }\] \[1-\cos 4x=\frac{ 1 }{ 2 },\cos 4x=\frac{ 1 }{ 2 }=\cos \frac{ \pi }{3 },\cos \left( \frac{ -\pi}{ 3 } \right)\] \[4x=2k \pi+\frac{ \pi }{ 3 },2k \pi-\frac{ \pi }{ 3},k=0,1,2,3\] calculate x

OpenStudy (anonymous):

wow man thanks a lot!!

OpenStudy (anonymous):

yw